OpenStudy (anonymous):

3x – 2y = 8 x = 2 substitute the second part x=2 into the first part 3(2)-2y=8 6-2y=8 -2y=2 y=-1 then plug -1 in for y and solve for x 3x-2(-1)=8 3x+2=8 3x=6 x=2 i need the check step to this please.

OpenStudy (anonymous):

Yes looks right

OpenStudy (anonymous):

you should get x=2 back when you plug in y=-1

OpenStudy (anonymous):

3(2)-2(-1)=6+2=8 yes.

OpenStudy (anonymous):

sweet that what i thought th check step was but theres more to others peoples

OpenStudy (anonymous):

here is an example of how she wants it 3 + 2(9) = 21 3 + 18 = 21 21 = 21 Checks 9 = 3(3) 9 = 9 Check

OpenStudy (anonymous):

and i am not sure how to get that with my problem